What is the consumer surplus and producer surplus for : Qd= 300-3p and Qs= 50+2p


1
Expert's answer
2022-03-29T12:11:47-0400

"300-3p = 50+2p"

"250 = 5p"

"P_{e} = 50"

While "Q_{e} = 300- 3(50) = 150"

Intercept for demand curve;

"0 = 300-3P"

"P = 100"

Intercept for supply;

"0 = 50 + 2P"

"P = -25"


The producer surplus is as follows;

"\frac{1}{2}\times (50--25)\times 150 =5625"


The consumer surplus is;

"\frac{1}{2}\times 150\times 50 = 3750"